Estimator properties:

6 Estimators properties 6.1 Exercise 1 In order to estimate the average number of hours that children spend watching tv, a Bernoulli sample of size n = 5 children was selected from a primary school. Let X be the variable that represents the hours spent watching tv, let E(X)-μ the parameter to estimate and var(X-σ2 the variance. Compare the following two proposed estimators Τι 1. Compare the two estimators for u on the basis of their bias 2. Compare the two esimators for u on the basis of their efficiency 3. Which estimator is preferable for estimating u?6.2 Exercise 2 Let X1, X2, X3 be a bernoullian sample of size n-3 from a population X of var(X) σ2. Compare the following three estimators for the unknown mean of the population E(X-μ: 12 T2-(X1 + 2X2 +3X3) 3-10 I. Compare the two estimators for μ on the basis of their bias 2. Consider the unbiased estimators and evaluate their efficiency.6.3 Exercise 3 In order to make sure his emplyees do not work on weekends, the HR manager monitors the number of emails they received last weekend. The following data represent a Bernoulli sample of 7 employees. σ2 the variance, for Let E(X) estimating u he was suggested to use the following two estimators: μ the parameter to estimate and var(X)i=1 I. Find the estimate for μ using the two proposed 2. Are the two estimators unbiased? 3. Which of them is more efficient? estimators
Answer #1

Solution

Exercise 1

Back-up Theory

A statistic, T (from sample) is an unbiased estimator of a parameter (of population) θ, if

E(T) = θ………………………………………………………………………………… (1)

Bias of an estimator, B = E(T) - θ………………………………………………………(2)

E{∑(i = 1 to n)(aiXi)} = ∑(i = 1 to n){aiE(Xi)}………………………………………….. (3)

V{∑(i = 1 to n)(aiXi)} = ∑(i = 1 to n){ai2V(Xi)}, if Xi are independent………………. (4)

Now to work out the solution,

Given X1, X2, X3, X4, X5, are iid with mean µ and variance σ2,

Part (1)

Vide (3),

E(T1) = E{(1/5)∑(i = 1 to 5)(Xi)}

=(1/5) ∑(i = 1 to 5){E(Xi)}

= (1/5)∑(i = 1 to 5)µ

= µ

So, vide (1), T1 is an unbiased estimator of µ and hence vide (2), bias (T1) = 0.

Vide (3),

E(T2) = E{(1/5)(2X1 – X2 + 2X3 + X4 + X5}

= (1/5){2E(X1) – E(X2) + 2E(X3) + E(X4) + E(X5)}

= (1/5)(2µ - µ + 2µ + µ + µ)

= µ

So, vide (1), T2 is an unbiased estimator of µ and hence vide (2), bias (T2) = 0.

Thus, both estimators have equal bias of zero.Answer

Part (2)

Variance is a measure of efficiency of an estimator.

Vide (4),

V(T1) = V{(1/5)∑(i = 1 to 5)(Xi)}

= (1/25)∑(i = 1 to 5)V(Xi)}

= (1/25)∑(i = 1 to 5)σ2

= (1/25)5σ2

= (1/5)σ2

V(T2) = V{(1/5)(2X1 – X2 + 2X3 + X4 + X5}

= (1/25){4V(X1) + V(X2) + 4V(X3) + V(X4) + V(X5)}

= (1/25)(4σ2 + σ2 + 4σ2 + σ2 + σ2)

= (11/25)σ2

Thus, V(T2) > V(T1) Answer

Part (3)

Since V(T1) < V(T2), T1 is more efficient that T2. Answer

DONE